OpenStudy (hitaro9):

Evaluate the Legendre Symbol. (19/67)

OpenStudy (hitaro9):

I could use Euler's Criterion here, but a lot of the previous problems have been relatively easy applications of theorems or were easier to evaluate. This seems pretty unclean to solve. Is there a combination of two theorems I'm just not seeing?

ganeshie8 (ganeshie8):

studied quadratic reciprocity law yet ?

OpenStudy (hitaro9):

Next question had something on it, I hadn't really looked at it yet. It had some terms I wasn't familiar with

ganeshie8 (ganeshie8):

quadratic reciprocity law takes time to understand, it took me 2-3 weeks to properly understand how it works and everything

ganeshie8 (ganeshie8):

i said that so you won't approach it thinking it is just another theorem like euler's criterion

OpenStudy (hitaro9):

Alright then. I'm guessing there's no way to easily break down the problem without it then? I see this lemma involving sums in the book, is this what we should use? 

OpenStudy (hitaro9):

Where p =67 and a = 19

ganeshie8 (ganeshie8):

that looks like a step in the proof of quadratic reciprocity law

ganeshie8 (ganeshie8):

You can evaluate ANY legendre symbol in one line with easy arithmetic using quadratic reciprocity law

ganeshie8 (ganeshie8):

i don't want this to be your introduction to quadratic reciprocity law if you haven't studied it yet..

OpenStudy (hitaro9):

Oh wow okay. Would it be worth attempting to learn it now for this assignment or should I wait until I have more time to focus on it?

ganeshie8 (ganeshie8):

wait until your professor teaches you officially

OpenStudy (hitaro9):

Alright, so should I just go through with Euler's Criterion then? Even if it is a bit of a pain?

OpenStudy (hitaro9):

Or do you see any alternative theorems that could be applied?

ganeshie8 (ganeshie8):

With euler's citerion, you need to evaluate\[(19/67)=19^{(67-1)/2}\pmod{67}\]

OpenStudy (hitaro9):

Right, and those are some pretty big numbers but I can go through it like any other mod problem

ganeshie8 (ganeshie8):

yeah i see its not reducing nicely

OpenStudy (hitaro9):

Is the law of quadratic reciprocity something that can be easily applied if not understood?

ganeshie8 (ganeshie8):

notice that both 19 and 67 are of form 4k+3, so\[(19/67)=-(67/19)\]

ganeshie8 (ganeshie8):

pretend we can do that, can you evaluate that legendre symbol ?

ganeshie8 (ganeshie8):

-(67/19) should be relatively easy because the modulus is just 19 now

OpenStudy (hitaro9):

That looks a lot more manageable 1 moment

ganeshie8 (ganeshie8):

first reduce 67 in mod 19

OpenStudy (hitaro9):

Neither of which is too terrible

ganeshie8 (ganeshie8):

careful, don't forget the negative..

OpenStudy (hitaro9):

Right that goes to 10, which could either evailuate at 10^9 mod 19 or (2/19) (5/19) mod 19

OpenStudy (hitaro9):

Right

OpenStudy (hitaro9):

The negative applies to legendre value right? Not what's inside?

OpenStudy (hitaro9):

So we'll apply that at the end

ganeshie8 (ganeshie8):

\[(19/67)=-(67/19)=-(10/19)=-(2/19)(5/19)\]

ganeshie8 (ganeshie8):

okay looks good!

ganeshie8 (ganeshie8):

with quadratic reciprocity law, you don't really need to work 2^9 and 5^9

ganeshie8 (ganeshie8):

apply quadratic reciprocity law for (5/19) again

ganeshie8 (ganeshie8):

since 5 is of form 4k+1 : \[(5/19) = (19/5)\]

OpenStudy (hitaro9):

Oh okay then, neat. This is pretty useful, thank you.

ganeshie8 (ganeshie8):

\[(19/67)=-(67/19)=-(10/19)=-(2/19)(5/19)\\=-(2/19)(19/5)=-(2/19)(-1/5)\\=-(-1)(1)\\=1\]

ganeshie8 (ganeshie8):

basically any legendre symbol can be reduced to finding the quadratic residue of one of the integers less than or equal to 3

OpenStudy (hitaro9):

Yeah, still working through the arithmetic, but I can definitely see how this can help reduce large numbers

OpenStudy (hitaro9):

Thank you so much for all the help you've been tonight (and previous nights as well haha)

ganeshie8 (ganeshie8):

np